# Test-time Correlation Alignment

### v1

@@ 5.1 Experimental settings
We evaluate the adaptation performance on two main tasks: image domain adaptation and image corruption adaptation. Following previous studies, for domain adaptation, we use the PACS (Li et al., 2017) dataset and the OfficeHome (Venkateswara et al., 2017) dataset. For image corruption adaptation, we utilize the CIFAR-10C and CIFAR-100C (Hendrycks & Dietterich, 2019) datasets. The comparison methods we employ include: BN (Schneider et al., 2020), TENT (Wang et al., 2020) EATA (Niu et al., 2022), SAR (Niu et al., 2023), TSD (Wang et al., 2023), TIPI (Nguyen et al., 2023), and TEA (Yuan et al., 2024). Backbone networks include ResNet-18/50 (He et al., 2016) and ViT-B/16 (Dosovitskiy, 2020). Additionally, the evaluation encompasses multiple performance aspects, including accuracy, efficiency, and resistance to forgetting. For LinearTCA+, we report its results combined with the best baseline. Refer to Appendix D for more implement information. For further experimental results and analysis, please see Appendix E.
@@ 5.2 Experimental validation of theories
For Theorem 3.5: Correlation of high-certainty test instances approximates the source correlation. We divide the test embeddings of CIFAR-10C under ResNet-18 into 10 groups based on prediction uncertainty and calculate the correlation distance between each group and the original source. As shown in Figure 3a, groups with lower uncertainty exhibit smaller correlation distances, indicating a closer approximation to the source correlation.
@@ 5.3 Comparison with TTA Methods
Accuracy. Table 1 presents ACC comparisons between TCA methods and state-of-the-art TTA approaches across various benchmarks, backbones, and tasks. (1) As a plug-and-play module, LinearTCA+ consistently enhances performance across all datasets and backbones, achieving a new state-of-the-art. Notably, on the CIFAR-10C dataset with the ViT-B/16 backbone, LinearTCA+ shows substantial improvements over the best-performing baseline, with an increase of 5.88%. (2) Across datasets, LinearTCA shows robust improvement compared to the source model, with average gains of 1.79%, 1.69%, 8.48%, and 3.67%, respectively. Particularly, on the OfficeHome dataset, LinearTCA consistently outperforms most baseline methods. However, on datasets such as CIFAR-10/100C, although LinearTCA yields ACC gains of 8.48% and 3.67% over the source model, it falls short of some advanced methods. (3) Across backbones, LinearTCA also demonstrates robust improvements compared to the source model, especially with the ViT-B/16 backbone, surpassing the highest-performing baseline on most datasets. We provide a detailed analysis of these experimental results in Section 5.4 to further reveal the strengths and limitations of LinearTCA.
Method Memory(MB)
ResNet-18 ResNet-50 ViT-B/16 AVG
SOURCE 920.61 878.87 917.02 905.50
BN +0.25 +48.57 - -
SAR +2642.82 +5380.18 +5401.31 +4474.77
EATA +5332.44 +10838.33 +11175.83 +9115.53
TENT +1883.63 +4788.93 +5246.53 +3973.03
TSD +2023.27 +5156.26 +9280.50 +5486.68
TEA +7316.95 +15735.97 +16082.00 +13044.97
TIPI +2520.01 +10660.83 +12542.71 +8574.52
@@
Method Time(s)
ResNet-18 ResNet-50 ViT-B/16 AVG
SOURCE 13.40 20.86 21.00 18.42
BN +0.06 +3.99 - -
SAR +14.06 +31.93 +63.12 +36.37
EATA +0.56 +9.07 +16.52 +8.72
TENT +5.59 +21.50 +37.26 +21.45
TSD +7.38 +17.72 +32.41 +19.17
TEA +130.18 +302.07 +627.85 +353.37
TIPI +31.65 +62.73 +76.35 +56.91
[…]
Table 2: Maximum GPU memory usage and running time of different TTA methods on CIFAR-10C.
@@
Table 3: Independent maximum GPU memory usage of LinearTCA on CIFAR-10C.
Efficiency. We assess the efficiency of each method from two aspects: maximum GPU memory usage and total runtime. Table 2 presents the experimental results for each method on the CIFAR-10C dataset with different backbones. Our TCA method consistently achieves the lowest memory and time consumption across all backbones. In terms of memory usage, since we record peak memory consumption, LinearTCA exhibits minimal independent memory usage (as shown in Table 3) and thus does not impose additional memory constraints on the device. In contrast, other methods are embedded within the model’s forward and backward propagation processes, significantly increasing peak memory usage (e.g., TEA’s maximum memory usage is 14 times that of Source). Regarding runtime, when the backbone is ViT-B/16, LinearTCA’s time consumption is on average only 6 of the best baseline EATA. These results demonstrate LinearTCA’s exceptional efficiency, making it particularly suitable for deployment on resource-constrained edge devices.
Forgetting resistance. Table 4 presents the changes in ACC when each method, with ResNet-18 as the backbone, returns to the source domain after adaptation on various datasets. “LinearTCA w/o $W$” refers to the result obtained by directly removing the linear transformation $W$, which is entirely equivalent to source and does not lose any source domain information. Notably, even after applying the linear transformation, LinearTCA exhibits significantly better forgetting resistance compared to other methods. This is especially evident on the PACS dataset, where LinearTCA shows a “positive backward transfer” ability that even improves performance on the source domain. Additionally, LinearTCA+ significantly enhances the resilience to forgetting of other methods.
Method PACS OfficeHome CIFAR-10C CIFAR-100C AVG
SOURCE 99.35 94.40 92.36 70.39 89.12
BN 98.90 (-0.44) 92.85 (-1.55) 62.96 (-29.40) 37.63 (-32.76) 73.09 (-16.04)
@@
Table 4: The accuracy of different TTA methods when returning to the source domain after adaptation.
5.4 Analysis.
Effective range of LinearTCA. Notably, as discussed in Section 5.3, although LinearTCA+ significantly improves all TTA methods, LinearTCA only achieves SOTA performance on part of datasets and backbones. The reasons may be: 1) Although the highest-certainty embeddings are selected as pseudo-source domains, if these embeddings still exhibit substantial differences from the true source domain (or if the backbone’s feature extraction capacity is insufficient, e.g., ResNet-18 vs. ViT-B/16), the performance ceiling of LinearTCA is limited. In contrast, other TTA methods update the model, thereby raising this ceiling and facilitating easier correlation alignment for LinearTCA+. 2) We only use a linear transformation $W$ for alignment, which may work well for simple shifts; however, the true distribution shifts may not conform to linear transformations but exhibit complex nonlinear relationships. We design a demo experiment to validate this hypothesis. In Figure 4a and b, the test domain shifts are linear and nonlinear, respectively. As shown, the transformed embeddings in Figure 4a align well with the original distribution, while the performance in Figure 4b shows partial alignment which is still insufficient.
@@
Ablation study. Our method contains only one hyperparameter—the number of pseudo source domain embeddings, $k$. However, considering that the total number of test instances is unknown in practical applications, we also randomly sample $k_{2}$ embeddings from the overall test set to investigate the impact of $k_{2}$ on LinearTCA performance. As shown in Figure 4c and d, on the OfficeHome dataset, the accuracy of LinearTCA is highest when $k$ and $k_{2}$ are set to 10 and 2400, respectively. Notably, across a wide range of values, LinearTCA can perform better than source model, indicating that our method can be easily applied in practice.
@@ 6 Conclusion and Future Work
Future work may incorporate nonlinear transformations for more effective correlation alignment. Additionally, with the interesting “positive backward transfer” phenomenon observed in Table 4, we will further investigate the underlying mechanism.
@@ A.2 Test-Time Adaptation
Backpropagation (BP)-Free TTA: This group includes batch normalization (BN) calibration methods (Wu et al., 2024; Schneider et al., 2020) and certain pseudo-labeling methods (Zhang et al., 2023) that do not update model parameters. BN-based methods posit that the statistics in BN layers capture domain-specific knowledge. To mitigate the domain gap, these methods replace training BN statistics with updated statistics computed from the target domain. Some pseudo-labeling methods utilize prototype similarity or k-nearest neighbor (kNN) approaches to refine predictions. Although BP-Free TTA methods are computationally efficient, their domain adaptation capabilities are often limited.
Backpropagation (BP)-Based TTA: This group encompasses certain pseudo-labeling methods (Zeng et al., 2024), consistency training methods (Sinha et al., 2023), and clustering-based training methods (Lee et al., 2024). Some pseudo-labeling methods use filtering strategies, such as thresholding or entropy-based approaches, to generate reliable pseudo-labels, thereby reducing the discrepancy between predicted and pseudo-labels. For instance, TSD (Wang et al., 2023) filters unreliable features or predictions with high entropy, as lower entropy correlates with higher accuracy, and applies a consistency filter to refine instances further. Consistency training methods aim to enhance the stability of network predictions or features by addressing variations in input data, such as noise or perturbations, and changes in model parameters. TIPI (Nguyen et al., 2023), for example, simulates domain shifts via input transformations and employs regularizers to maintain model invariance. Clustering-based training methods leverage clustering techniques to group target features, and reduce uncertainty in predictions and improving model robustness. TENT (Wang et al., 2020) minimizes prediction entropy on target data, while EATA (Niu et al., 2022) selects reliable instances to minimize entropy loss and applies a Fisher regularizer. SAR (Niu et al., 2023) removes noisy instances with large gradients and encourages model weights to converge toward a flat minimum, enhancing robustness against residual noise. Generally, BP-Based TTA methods demonstrate superior domain adaptation capabilities compared to BP-Free methods, but they typically require multiple backward propagations for each test instance, leading to computational inefficiencies.

### v2

@@ 5.1 Experimental settings
Following previous studies, we evaluate the adaptation performance on two main tasks: domain generalization (PACS (Li et al., 2017), OfficeHome (Venkateswara et al., 2017), and DomainNet (Peng et al., 2019) dataset) and image corruption (CIFAR-10-C,CIFAR-100-C, and ImageNet-C (Hendrycks & Dietterich, 2019)). What’s more, we also evaluate our method on multimodal tasks based on CLIP (Radford et al., 2021). The comparison methods include backpropagation-free (BN (Schneider et al., 2020), T3A (Iwasawa & Matsuo, 2021), AdaNPC (Zhang et al., 2023)) and backpropagation-based methods (TENT (Wang et al., 2020), PLC (Lee, 2013), EATA (Niu et al., 2022), SAR (Niu et al., 2023), TSD (Wang et al., 2023), TIPI (Nguyen et al., 2023), TEA (Yuan et al., 2024)). Backbone networks include ResNet-18/50 (He et al., 2016) and ViT-B/16 (Dosovitskiy, 2020). Additionally, the evaluation encompasses multiple aspects, including accuracy, efficiency, and resistance to forgetting. For LinearTCA+, we report its results combined with the best baseline. Refer to Appendix D for more implement information. For further experimental results and analysis, please see Appendix E.
@@ 5.2 Experimental validation of theories
For Theorem 3.5: Correlation of high-certainty test instances approximates the source correlation. We divide the test embeddings of CIFAR-10-C under ResNet-18 into 10 groups based on prediction uncertainty and calculate the correlation distance between each group and the original source. As shown in Figure 3a, groups with lower uncertainty exhibit smaller correlation distances, indicating a closer approximation to the source correlation.
@@ 5.3 Comparison with TTA Methods
Accuracy. Table 1 presents ACC comparisons between TCA methods and state-of-the-art TTA approaches across various benchmarks, backbones, and tasks. (1) As a plug-and-play module, LinearTCA+ consistently enhances performance across all datasets and backbones, achieving a new state-of-the-art. Notably, on the CIFAR-10-C dataset with the ViT-B/16 backbone, LinearTCA+ shows substantial improvements over the best-performing baseline, with an increase of 4.95%. (2) Across datasets, LinearTCA shows robust improvement compared to the source model, with average gains of 1.79%, 1.75%, 1.78%, 8.48%, 3.67% and 4.51%, respectively. Particularly, on the OfficeHome and DomainNet dataset, LinearTCA outperforms all baseline methods. However, on datasets such as CIFAR-10/100-C and ImageNet-C, although LinearTCA yields ACC gains of 8.48%, 3.67% and 4.51% over the source model, it falls short of some advanced methods. (3) Across backbones, LinearTCA also demonstrates robust improvements compared to the source model, especially with the ViT-B/16 backbone, surpassing the highest-performing baseline on most datasets. We provide a detailed analysis of these experimental results in Section 5.5 to further reveal the strengths and limitations of LinearTCA.
Type Method Memory(MB) Time(s)
ResNet-18 ResNet-50 ViT-B/16 AVG ResNet-18 ResNet-50 ViT-B/16 AVG
SOURCE 920.61 878.87 917.02 905.50 3.92 9.16 3.98 5.69
BP-Free BN +0.25 +48.57 - - +0.88 +4.80 - -
T3A +1.00 +4.43 +2.02 +2.48 +1.98 +3.62 +12.22 +5.94
AdaNPC +2.04 +8.23 +2.96 +4.41 +1.73 +2.78 +12.08 +5.53
BP-Based TENT +1883.63 +4788.93 +5246.53 +3973.03 +3.85 +11.52 +12.27 +9.22
PLC +1934.14 +4787.26 +8624.95 +5115.45 +5.94 +9.51 +25.86 +13.77
EATA +5332.44 +10838.53 +11172.56 +9114.51 +1.76 +4.20 +22.81 +9.59
SAR +2642.82 +5380.18 +5401.31 +4474.77 +11.23 +23.31 +54.08 +29.54
TSD +2025.07 +5162.55 +9280.69 +5489.44 +4.70 +13.47 +34.68 +17.62
TEA +7316.95 +15733.10 +16082.00 +13044.02 +123.14 +278.87 +596.28 +332.76
TIPI +2520.01 +10660.83 +12542.71 +8574.52 +26.54 +49.73 +45.25 +40.51
@@
Table 2: Maximum GPU memory usage and running time of different TTA methods on CIFAR-10-C.
Efficiency. We evaluate each method’s efficiency in terms of peak GPU memory usage and total runtime. Table 2 reports results on the CIFAR-10-C dataset across different backbones. TCA consistently achieves the lowest memory and time cost. For memory, since we record peak memory consumption, LinearTCA exhibits minimal independent memory usage (as shown in Table 3) and thus does not impose additional memory constraints on the device.
@@
Table 3: Independent maximum GPU memory usage of LinearTCA on CIFAR-10-C.
In contrast, other methods are embedded within the model’s forward and backward propagation processes, significantly increasing peak memory usage (e.g., TEA uses 15$\times$ the memory of the source model). For runtime, with a ViT-B/16 backbone, LinearTCA requires only 0.6% of AdaNPC’s time. These results highlight LinearTCA’s high efficiency, making it well-suited for resource-constrained edge deployment.
Forgetting resistance. Table 4 shows the change in accuracy when each method (using ResNet-18) returns to the source domain after adaptation. “LinearTCA w/o $W$” refers to the variant without the linear transformation, which is equivalent to the original source model and thus retains full source knowledge. Despite applying $W$, LinearTCA demonstrates much stronger resistance to forgetting than other methods—especially on PACS, where it even improves source performance, showing positive backward transfer. Moreover, LinearTCA+ further enhances the forgetting robustness of existing TTA methods.
Type Method PACS OfficeHome CIFAR-10-C CIFAR-100-C AVG
BP-Free SOURCE 99.35 94.40 92.36 70.39 89.12
BN 98.90 (-0.44) 92.85 (-1.55) 62.98 (-29.38) 39.45 (-30.94) 73.55 (-15.58)
T3A 99.33 (-0.01) 93.31 (-1.09) 91.95 (-0.41) 65.66 (-4.73) 87.56 (-1.56)
AdaNPC 99.28 (-0.06) 93.31 (-1.09) 92.00 (-0.36) 63.88 (-6.51) 87.12 (-2.01)
[…]
BP-Based TENT 96.74 (-2.61) 92.79 (-1.61) 90.26 (-2.10) 67.27 (-3.12) 86.76 (-2.36)
PLC 97.12 (-2.23) 92.73 (-1.67) 63.05 (-29.31) 39.48 (-30.91) 73.09 (-16.03)
@@
Table 4: The accuracy of different TTA methods when returning to the source domain after adaptation. “BP-Free” indicates backpropagation-free TTA methods, while “BP-Based” denotes backpropagation-dependent ones.
5.4 Performance on Closed-Source Foundation Models
To validate TCA’s effectiveness on closed-source foundation models, we conduct experiments with CLIP (Radford et al., 2021) on PACS, OfficeHome, and VLCS datasets, following the experimental setup in WATT (Osowiechi et al., 2024). As shown in Table 5, TCA achieving performance improvements of 1.28%, 2.08%, and 2.85% on the three datasets respectively. The superior results stem from our method’s explicit alignment of embedding distributions with the source domain, which proves particularly effective for multi-modal models like CLIP that compute image-text similarity directly. While LinearTCA+ holds a slight advantage, both variants perform similarly, suggesting that even simple correlation alignment can notably enhance performance on popular models like CLIP. This underscores its effectiveness as a versatile plug-and-play module for improving diverse adaptation methods.
Method PACS AVG
A C P S
CLIP† 97.44 97.38 99.58 86.06 95.12
TENT† 97.54$\pm$0.02 97.37$\pm$0.04 99.58$\pm$0.00 86.37$\pm$0.05 95.22
TPT† 95.10$\pm$0.41 91.42$\pm$0.22 98.56$\pm$0.40 87.23$\pm$0.06 93.08
CLIPArTT† 97.64$\pm$0.02 97.37$\pm$0.02 99.58$\pm$0.00 86.79$\pm$0.04 95.35
WATT-P† 97.49$\pm$0.08 97.47$\pm$0.04 99.58$\pm$0.00 89.73$\pm$0.16 96.07
WATT-S† 97.66$\pm$0.08 97.51$\pm$0.02 99.58$\pm$0.00 89.56$\pm$0.14 96.08
@@ A C P S
Method OfficeHome AVG
A C P R
CLIP† 79.30 65.15 87.34 89.31 80.28
TENT† 79.26$\pm$0.14 65.64$\pm$0.05 87.49$\pm$0.02 89.50$\pm$0.04 80.47
TPT† 81.97$\pm$0.17 67.01$\pm$0.21 89.00$\pm$0.06 89.66$\pm$0.06 81.91
CLIPArTT† 79.34$\pm$0.05 65.69$\pm$0.11 87.35$\pm$0.07 89.29$\pm$0.03 80.42
WATT-P† 80.37$\pm$0.25 68.59$\pm$0.13 88.15$\pm$0.07 90.18$\pm$0.03 81.82
WATT-S† 80.43$\pm$0.09 68.26$\pm$0.11 88.02$\pm$0.08 90.14$\pm$0.06 81.71
[…]
Method VLCS AVG
C L S V
CLIP† 99.43 67.75 71.74 84.90 80.96
TENT† 99.43$\pm$0.00 67.31$\pm$0.14 71.57$\pm$0.15 85.10$\pm$0.11 80.85
TPT† 97.62$\pm$0.12 49.77$\pm$0.03 71.56$\pm$0.86 71.17$\pm$0.70 72.53
CLIPArTT† 99.43$\pm$0.00 67.74$\pm$0.10 71.67$\pm$0.01 84.73$\pm$0.08 80.89
WATT-P† 99.36$\pm$0.00 67.55$\pm$0.39 74.75$\pm$0.07 82.53$\pm$0.10 81.05
WATT-S† 99.36$\pm$0.00 68.59$\pm$0.25 75.16$\pm$0.12 83.24$\pm$0.05 81.59
[…]
Table 5: The accuracy comparison of different methods on PACS, OfficeHome, and VLCS datasets using CLIP-ViT-B/16. †: numbers are from WATT (Osowiechi et al., 2024). The best results are highlighted in boldface, and the second ones are underlined.
5.5 Analysis
Effective range of LinearTCA. As discussed in Section 5.3, although LinearTCA+ significantly improves all TTA methods, LinearTCA only achieves SOTA performance on part of datasets and backbones. The reasons may be: 1) Although the highest-certainty embeddings are selected as pseudo-source domains, if these embeddings still exhibit substantial differences from the true source domain (or if the backbone’s feature extraction capacity is insufficient, e.g., ResNet-18 vs. ViT-B/16), the performance ceiling of LinearTCA is limited. In contrast, other TTA methods update the model, thereby raising this ceiling and facilitating easier correlation alignment for LinearTCA+. 2) We only use a linear transformation $W$ for alignment, which may work well for simple shifts; however, the true distribution shifts may not conform to linear transformations but exhibit complex nonlinear relationships. We design a demo experiment to validate this hypothesis. In Figure 4a and b, the test domain shifts are linear and nonlinear, respectively. As shown, the transformed embeddings in Figure 4a align well with the original distribution, while the performance in Figure 4b shows partial alignment which is still insufficient. We further explore the utilization of nonlinear architecture (MLP) for calculating transformation $W$. As shown in Table 6, incorporating nonlinear activations with deeper architectures leads to further improvements.
@@ 5.5 Analysis
Backbone Method PACS AVG OfficeHome AVG
A C P S A C P R
ResNet-18 Source 78.37 77.39 95.03 76.58 81.84 56.45 48.02 71.34 72.23 62.01
LinearTCA 80.91 81.02 95.69 76.74 83.59 59.46 50.40 72.02 72.78 63.66
LinearTCA+ 88.38 87.12 96.59 83.00 88.77 59.83 51.80 72.29 73.17 64.27
LinearTCA(MLP-2) 81.24 81.73 95.89 78.15 84.25 59.62 50.84 72.07 72.94 63.87
LinearTCA+(MLP-2) 88.68 87.15 96.68 83.19 88.93 59.83 51.80 72.79 73.46 64.47
LinearTCA(MLP-3) 81.62 81.81 96.03 79.35 84.70 59.62 50.65 72.07 73.02 63.84
LinearTCA+(MLP-3) 88.38 87.23 96.59 83.36 88.98 59.83 52.08 72.79 73.54 64.56
Table 6: Extending LinearTCA/LinearTCA+ by introducing MLP-based transformations with two (MLP-2) and three (MLP-3) layers The best results are highlighted in boldface, and the second ones are underlined.
Ablation study. Our method involves only one hyperparameter—the number of pseudo-source embeddings $k$. Since the total number of test samples is often unknown in practice, we also sample $k_{2}$ embeddings from the test set to study its impact. As shown in Figure 4c,d, LinearTCA achieves the best accuracy on OfficeHome when $k=10$ and $k_{2}=2400$. Importantly, it consistently outperforms the source model across a wide range of $k$ and $k_{2}$, demonstrating strong practical applicability.
Backbone Method PACS AVG OfficeHome AVG
A C P S A C P R
ResNet-18 Source 78.37 77.39 95.03 76.58 81.84 56.45 48.02 71.34 72.23 62.01
LinearTCA 80.91 81.02 95.69 76.74 83.59 59.04 49.97 71.77 72.89 63.42
LinearTCA+ 88.38 87.12 96.59 83.00 88.77 59.33 51.18 72.20 71.72 63.61
TCA(a) 86.18 82.67 95.03 80.81 86.17 58.69 50.80 72.04 72.92 63.61
LinearTCA(b) 81.59 81.48 96.05 77.51 84.15 59.94 51.63 72.36 73.48 64.35
LinearTCA+(b) 88.98 87.57 96.74 83.30 89.15 60.03 52.29 72.55 73.87 64.64
Table 7: Upper performance bound for TCA. TCA(a): Fine-tuning directly on the target distribution. LinearTCA(b) and LinearTCA+(b): Applying LinearTCA and LinearTCA+ with real source distributions.
Method Art Domian of OfficeHome AVG
Batch Size 1 2 4 8 16 32 64 128 256 512 1024
Estimation error 2542 2414 2434 2430 2415 2417 2437 2415 2413 2424 2427 2433
Source 56.45 56.45 56.45 56.45 56.45 56.45 56.45 56.45 56.45 56.45 56.45 56.45
TEA 0.824 18.01 40.79 49.23 55.54 55.71 57.35 58.55 57.11 57.82 57.93 46.26
[…]
Table 8: Accuracy comparisons of different TTA methods on the Art domain of OfficeHome dataset with varying batch sizes based on ResNet-18. The best results are highlighted in boldface, and the second ones are underlined.
Upper performance bound for TCA. To assess the upper bound of TCA, we conduct two additional experiments in Table 8: (a) fine-tuning directly on the target domain; (b) applying LinearTCA and LinearTCA+ with real source distributions. Compared to the original LinearTCA+, approach (b) further improves performance, by 0.38% on PACS and 1.03% on OfficeHome. Both (a) and (b) outperform the original LinearTCA in most domains. On OfficeHome, even the simpler LinearTCA with real source data (b) surpasses fine-tuning (a), highlighting the importance of source distribution and the effectiveness of approximating it in TCA.
Performance under difference batch sizes. To study the impact of batch size, we evaluate TCA’s performance and pseudo-source estimation error under varying batch sizes in Table 8. Even with batch size 1, LinearTCA outperforms the source model by 2.16%, and LinearTCA+ consistently improves over TEA across all settings. This robustness stems from TCA’s incremental estimation of test-domain covariance, which converges over time. While small batch sizes mainly affect early predictions, their influence diminishes as more data is seen. Moreover, the pseudo-source estimation error remains unaffected by batch size, since it relies on a small set of high-confidence samples (Figure 4c) and benefits from the same incremental computation.
@@ 6 Conclusion and Future Work
Future work may incorporate more nonlinear transformations for more effective TCA. Additionally, with the interesting “positive backward transfer” phenomenon in Table 4, we will further investigate the underlying mechanism.
@@ A.2 Test-Time Adaptation
Backpropagation (BP)-Free TTA: This group includes batch normalization (BN) calibration methods (Wu et al., 2024; Schneider et al., 2020) and certain pseudo-labeling methods (Zhang et al., 2023) that do not update model parameters. BN-based methods posit that the statistics in BN layers capture domain-specific knowledge. To mitigate the domain gap, these methods replace training BN statistics with updated statistics computed from the target domain. Some pseudo-labeling methods such as T3A (Iwasawa & Matsuo, 2021) utilize prototype similarity and AdaNPC (Zhang et al., 2023) utilize k-nearest neighbor (kNN) to refine predictions. Although BP-Free TTA methods are computationally efficient, their image corruption adaptation capabilities are often limited.
Backpropagation (BP)-Based TTA: This group encompasses certain pseudo-labeling methods (Zeng et al., 2024), consistency training methods (Sinha et al., 2023), and clustering-based training methods (Lee et al., 2024). Some pseudo-labeling methods use filtering strategies, such as thresholding or entropy-based approaches, to generate reliable pseudo-labels, thereby reducing the discrepancy between predicted and pseudo-labels. For instance, PLC (Lee, 2013) updates classifier layer parameters with certain pseudo-labels during adaptation. TSD (Wang et al., 2023) filters unreliable features or predictions with high entropy, as lower entropy correlates with higher accuracy, and applies a consistency filter to refine instances further. Consistency training methods aim to enhance the stability of network predictions or features by addressing variations in input data, such as noise or perturbations, and changes in model parameters. TIPI (Nguyen et al., 2023), for example, simulates domain shifts via input transformations and employs regularizers to maintain model invariance. Clustering-based training methods leverage clustering techniques to group target features, and reduce uncertainty in predictions and improving model robustness. TENT (Wang et al., 2020) minimizes prediction entropy on target data, while EATA (Niu et al., 2022) selects reliable instances to minimize entropy loss and applies a Fisher regularizer. SAR (Niu et al., 2023) removes noisy instances with large gradients and encourages model weights to converge toward a flat minimum, enhancing robustness against residual noise. Generally, BP-Based TTA methods demonstrate superior domain adaptation capabilities compared to BP-Free methods, but they typically require multiple backward propagations for each test instance, leading to computational inefficiencies.